OpenStudy (anonymous):

The very rigorous medical program at a local university has a 25% drop out rate for each year. If the school admits 1,000 freshmen, how many diplomas will need to be ordered 4 years later?

OpenStudy (anonymous):

So each year, a quarter of 1000 freshmen will drop out. So: Year 1: \(\ \huge 1000 \times .25 = 250, so 750 remain \). Year 2: \(\ \huge 750 \times .25 \approx 188; so 562 remain \). Year 3: \(\ \huge 188 \times .25 = 47, so 141 remain \). Graduation Year: \(\ \huge 141 \times .25 \approx 35, so 106 remain \). So, the school will need \(\ \huge 106 \color{green} diplomas for this class \).

OpenStudy (anonymous):

Thats not anoption, its either 656,410,316,240

OpenStudy (anonymous):

\(\ \huge hmm.... \).

OpenStudy (anonymous):

it was 316

OpenStudy (zarkon):

but do you understand why?

OpenStudy (zarkon):

\[1000\times(.75)^4\]

OpenStudy (anonymous):

Yes because you subtract 25% of the students each year

OpenStudy (zarkon):

or you keep 75% each year